OpenStudy (chosenmatt):

Which equation best represents the situation? A teacher prepared math problems for students to solve in groups. She split her class into 4 groups. Each group got an equal number of problems. After distributing them, there were 2 left over. Group A completed their problems plus 2 extras (30 problems). How many problems did the teacher originally prepare?

OpenStudy (chosenmatt):

isit 4x+2=30?

OpenStudy (aripotta):

30 is the amount Group A completed, not the total number of problems for all groups

OpenStudy (aripotta):

do you have choices?

OpenStudy (chosenmatt):

yes

OpenStudy (chosenmatt):

x/4 +2 =30, x/2-4=30, 2x-4=30 and the one i said i thought it was

OpenStudy (aripotta):

the first one makes more sense

OpenStudy (chosenmatt):

x/4 +2 =30 this one? @AriPotta

OpenStudy (aripotta):

yea

OpenStudy (chosenmatt):

i got 1 more @AriPotta

OpenStudy (chosenmatt):

Which inequality best represents the situation? A pine tree grows at a rate of 1.5 feet per year. James plants a 2-foot tall pine tree in his yard. How many years will it take for the tree to be at least 14 feet tall? A. 2x + 1.5 ≥ 14 B. 2x – 1.5 ≤ 14 C. 1.5x + 2 ≥ 14 D. 1.5x – 2 ≤ 14 @AriPotta

OpenStudy (aripotta):

so x is the amount of years. 1.5 is the ft per yr. so 1.5x would be the amount of ft the tree grew in x yrs. once we add 2, it'll give us the amount of ft tall it is after x yrs. so 1.5x + 2 and "at least 14 ft tall" means greater than or equal to. so \[1.5x + 2\ge14\]

OpenStudy (aripotta):

did that make sense?

OpenStudy (chosenmatt):

yup

OpenStudy (aripotta):

cool

OpenStudy (chosenmatt):

You want to know more about Dante's homework. Which situation about his homework best represents the inequality 3x + 12 ≤ 48? A. Dante spends 12 minutes reading the lesson and 3 minutes completing each math problem. How many problems can he finish in 48 minutes? B. Dante spends 12 minutes completing each math problem and 3 minutes reading the lesson. How many problems can he finish in 48 minutes? C. Dante spends 12 minutes on his math homework and 3 minutes to complete each problem. How long does it take him to finish 48 problems? D. Dante completes 12 math problems and reads 3 pages in his math book. How many problems can he finish in 48 minutes? @AriPotta

OpenStudy (aripotta):

i think a